OpenStudy (anonymous):

hey can anybody tell me how to solve the improper fraction (8-4x+11x^2-2X^3)/(1-(1/4)X)(1-x+(1/2)x^2) using partial fractions

OpenStudy (anonymous):

if i were you i would multiply top and bottom by 4 to clear those stupid fraction before i began

OpenStudy (anonymous):

\[\frac{(8-4x+11x^2-2x^3)}{(1-\frac{1}{4}x)(1-x+\frac{1}{2}x^2)}\] is that it?

OpenStudy (anonymous):

actually i wold multiply top and bottom by 8. i hate working with fractions

OpenStudy (anonymous):

ya right

OpenStudy (anonymous):

\[\frac{64-32x+88x^2-16x^3}{(4-x)(2-2x-x^2)}\] yes?

OpenStudy (anonymous):

no that is not right

OpenStudy (anonymous):

\[\frac{64−32x+88x^2−16x^3}{(4−x)(2−2x+x^2)}\]

OpenStudy (anonymous):

but this is not an improper fraction because numerator and denominator both have degree 3

OpenStudy (anonymous):

whenever degrees r equal it is said to b improper fraction

OpenStudy (anonymous):

oh crap, i guess you can do it, but it will be a real pain. denominators will be \[2-2x+x^2\] and \[(4-x)^2\]

OpenStudy (anonymous):

u just tell me which method i shd use here without multiplying 8 to the numerator....bcoz i wanna retain those fracttions of denominator as it is

OpenStudy (anonymous):

oh maybe not. ok we can start. first write \[\frac{A}{1-\frac{1}{4}x}+\frac{Bx+C}{1-x+\frac{1}{2}x^2}=\frac{8-4x+11x^2-2x^3}{(1+\frac{1}{4}x)(1-x-\frac{1}{2}x^2)}\]

OpenStudy (anonymous):

now cross multiply on the left to add the fractions. you will have \[A(1-x+\frac{1}{2}x^2)+(Bx+C)(1-\frac{1}{4}x)=x-4x-=+11x^2-2x^3\]

OpenStudy (anonymous):

ya k fine got it ....thank u

OpenStudy (anonymous):

typo on that line. the last = sign should not be there

OpenStudy (anonymous):

then multiply out, and combine like terms. then equation like coefficients. it is a pain

OpenStudy (anonymous):

ya k fine got it ....thank u

OpenStudy (anonymous):

yw